A vertical rod of length l is moved with constant velocity v towards east. The vertical component of earth magnetic field is B and angle of dip is theta. What is the induced emf in the rod?

Grade:12th pass

4 Answers

Ritik
5 years ago
Induced EMF is given by E= Blv ......1 Vertical component of B is B sin theta From 1 ,we getE= (B sin theta)× L× v cos theta E= B/v cot theta
1234567890
4 years ago
This petition is made in regards to Hip Hop international Artist Daniel Hernandez stage name (Tekashi 69) His Family , Fans , Friends & Supporters hope that the judge can grant this young man bail so he can get back with his family & continue to make great music for his fans while he fights this legal battle. Let’s show that this Rainbow hair rapper is a non threat to any community but is loved for all that he has done & the positive impact he has on others. We thank everyone for your love & support 
ankit singh
2 years ago
A vertical conducting rod of length l is moved at a constant velocity v eastward. The vertical component of the Earth's magnetic field is B and the angle of dip is theta. ... The rod is vertical and moves east ward, therefore it 'cuts' horizontal component Bh of the magnetic field.
Naman Agrawal
one year ago

From the above figure, are conclude that magnetic force which is responsible for induction of emf is horizontal component of magnetic field only.

 
Formula Δe=(Vˉ×Bˉ)l=VBlsin90=VBl
 
Now tanθ=BBHBH=Bcotθ
 
Now, emf induced = 
e=V(BHl)
 
e=VBlcotθ
 
e=Blvcotθ